OpenStudy (maheshmeghwal9):

Solve:

OpenStudy (maheshmeghwal9):

\[(4x+3/4)-(x-(2x-1/3)=x+(1/3).\]

OpenStudy (zepp):

What would happen if you try to remove the parentheses first?

OpenStudy (diyadiya):

\[\left( 4x+\frac{3}{4} \right)-\left( x-\left( 2x- \frac{1}{3} \right) \right)=x+ \frac{1}{3} ???\]

OpenStudy (maheshmeghwal9):

- becomes +

OpenStudy (zepp):

Yup, can you do it? :)

OpenStudy (maheshmeghwal9):

ya diya

OpenStudy (maheshmeghwal9):

Please solve!

OpenStudy (diyadiya):

remove the parentheses ...

OpenStudy (maheshmeghwal9):

ok!

OpenStudy (maheshmeghwal9):

I will do it now .Thanx to allll!:)

OpenStudy (zepp):

Good :)